A particle is fired at 22 degrees with respect to an incline and then hits a point on the incline as label. If the initial velocity is 15 m/s, a. find the time it takes for the projectile to hit the incline plane. b. find the distance from launching point to the landing point.


1
Expert's answer
2021-05-31T15:14:03-0400

Gives

"\theta=22°"

Velocity= 15 m/sec

Part (a)

Time take projectile to hit the inclined plane

"t=\frac{usin\theta}{g}"

"t=\frac{15\times sin22}{9.8}"

"t=0.574" sec

Total time launching point to landing point

"T=\frac{2usin\theta}{g}"

Put value

"T=\frac{2\times15\times sin22}{9.8}"

"T=1.15" sec

Part(b)

Distance from lonching to landing

"R=\frac{u^2sin2\theta}{g}"

Put value

"R=\frac{15^2\times sin44}{9.8}" =15.94m